OpenStudy (xapproachesinfinity):

convergence test! \[\sum_{1}^{\infty}\frac{n!e^n}{n^n}\] we are doing in the class power series and trying to test does the series \[\sum_{1}^{\infty}\frac{n!x^{2n}}{n^{n}} ~~for~~ x=\sqrt{e}\] converge or not

OpenStudy (xapproachesinfinity):

i tried to show that the limit of the nth term is not 0 therefore the series must diverge seems something is not good

OpenStudy (rational):

everything looks good, how did you compute the limit ?

OpenStudy (xapproachesinfinity):

hmm this is what i did \[\lim_{n\to \infty }n!(\frac{e}{n})^n=L, ~~L\ne0\] applied log to both sides but seems i have to difficulty to go further lol

OpenStudy (rational):

try using stirling approximation http://mathworld.wolfram.com/StirlingsApproximation.html

OpenStudy (xapproachesinfinity):

hmm ok let me see

OpenStudy (xapproachesinfinity):

\[\lim n\ln-n+n-n\ln n=\ln L \Longrightarrow L=1 \]

OpenStudy (xapproachesinfinity):

Am i right? using that approach lnn! is nlnn-n

OpenStudy (xapproachesinfinity):

seems that the limit on the left side is 0

OpenStudy (rational):

let me get paper and pen

OpenStudy (xapproachesinfinity):

alright :)

OpenStudy (rational):

we may use this directly instead http://gyazo.com/8165782b98540ff87489e00fa793bf97

OpenStudy (rational):

replace \(n!\) by \(n^ne^{-n}\sqrt{2\pi n}\)

OpenStudy (rational):

\[\lim\limits_{n\to\infty}\frac{\color{blue}{n!}e^n}{n^n}=\lim\limits_{n\to\infty}\frac{\color{blue}{n^ne^{-n}\sqrt{2\pi n}}e^n}{n^n}\] simplify and evaluate

OpenStudy (xapproachesinfinity):

that goes to infinity which is good

OpenStudy (xapproachesinfinity):

thanks :)